OpenStudy (anonymous):

What is the slope of the line passing through the points (6, -4) and (8, -7)? A. -3/2 B. -2/3 C. -11/14 D. 11/2

OpenStudy (nerdgirl):

I won't do it for you but I can help you solve it ok?

OpenStudy (anonymous):

Yeahhh

OpenStudy (nerdgirl):

Ok, so the formula for slope is Y2 - Y1 / X2 - X1. substitute those numbers in to that equation

OpenStudy (anonymous):

just doing y2-y1 u can figure it out

OpenStudy (nerdgirl):

Did you get -3/2?

OpenStudy (anonymous):

what is -7+4

OpenStudy (anonymous):

if u got -3/2 ur right

OpenStudy (anonymous):

-3

OpenStudy (nerdgirl):

You should always do it all the way through and never cut corners I've done that too many times and failed on tests. Ok?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

What is the x-intercept of the following line? 4x -7y = -28 A. (4, 0) B. (7, 0) C. (-4, 0) D. (-7, 0)

OpenStudy (nerdgirl):

Even if when it gives you the first number and you know the answer - always do it all the way through, otherwise in the future you won't know how to do it.

OpenStudy (nerdgirl):

Put that equation in y = mx+b form. First, subtract 4x from 28, so it looks like -7y = -28 - 4x. Ok?

OpenStudy (nerdgirl):

Oh wait my mistake sorry!! don't do that instead make it look like -7y = -4x - 28. Once you've done that, divide both sides by 7.

OpenStudy (anonymous):

yea I remember doing that in class